OpenStudy (anonymous):

If A is (1,2), B is (1,-6), and C is (4,7), what is the area of the triangle ABC?

hero (hero):

Find the lengths of segments AB, BC, and CA then use Heron's Formula.

OpenStudy (anonymous):

Could you please show me how to do that?

hero (hero):

You have to first plot the points on an xy plane. It is best for you to familiarize yourself with a graphing program like http://www.desmos.com for plotting points, segments and lines.

OpenStudy (anonymous):

Ok, then how do you find the length of the segments?

hero (hero):

Hang on a minute....

hero (hero):

Here's the triangle https://www.desmos.com/calculator/w6m2pzx6a2

hero (hero):

You can use distance formula to find the lengths of the segments.

hero (hero):

If you click "Projector Mode" you would be better able to keep track of the units.

hero (hero):

Obviously since the blue segment is completely vertical, it is easy to figure out its unit length of 8 units.

hero (hero):

The unit length of the green segment is, \[\sqrt{5^2 + 3^2} = \sqrt{25 + 9} = \sqrt{34}\]

hero (hero):

The unit length of the purple segment is \[\sqrt{13^2 + 3^2} = \sqrt{169 + 9} = \sqrt{178}\]

hero (hero):

So now you have triangle segment lengths of \(8, \sqrt{34}, \sqrt{178}\)

hero (hero):

Now use Heron's Formula to find the Area of the triangle.